# Euler-type approximation for the invariant measure: An abstract framework

### v1

@@ 2.3. Probabilistic representation
$W_{p}^{p}(\mu^{1},\mu^{2})=$ $\mathbb{E}(d^{p}(X^{1},X^{2})),$ (2.21)
$\mathbb{E}(d^{p}(\mathcal{X}_{s,t}^{1},\mathcal{X}_{s,t}^{2}))\leq$ $(1-b_{\ast}(t-s))\mathbb{E}(d^{p}(X^{1},X^{2}))+C_{\ast}\Gamma_{p}(\mu^{1},\mu^{2})(t-s)^{1+\varepsilon},$ (2.22)
@@
$W_{p}^{p}(\Theta_{s,t}^{1}(\mu^{1}),\Theta_{s,t}^{2}(\mu^{2}))$ $\leq$ $\mathbb{E}(d^{p}(\mathcal{X}_{s,t}^{1},\mathcal{X}_{s,t}^{2}))$
$\leq$ $(1-b_{\ast}(t-s))\mathbb{E}(d^{p}(X^{1},X^{2}))+C_{\ast}\Gamma_{p}(\mu^{1},\mu^{2})(t-s)^{1+\varepsilon}$
@@
$M_{p}(\Theta,\mu)\leq\left\|\mu\right\|_{p}^{p}+C\frac{e^{b}}{b}.$ (2.23)
@@ 3.1. Using the sewing lemma to prove coupling
We thus get that $\Theta^{1}$ and $\Theta^{2}$ are $(1,b_{\ast},\varepsilon)$-coupled. And by Lemma 2.6, $\Theta^{2}$ is also $1$-bounded. ∎
@@
  • If we are able to prove independently (relying on continuous time arguments) that $\theta$ is $1$-bounded and is $(1,b_{\ast},\varepsilon)$ self-coupled, then $\Theta$ and $\theta$ are $(1,b_{\ast},\varepsilon)$-coupled and $\Theta$ is $1$-bounded. In particular, if $\theta$ is time homogeneous and $\nu$ is the invariant measure for $\theta$, and if we take the time sequence $t_{k}=s+\sum_{i=1}^{k}\frac{1}{i+h^{-1}}$ with $h\in(0,1)$ as in Theorem 2.12 then, by (2.20), for every $\mu\in\mathcal{P}_{1}(B)$,
@@ 3.2. A Foster-Lyapunov type criterion to prove $p$-boundedness
Suppose that the family $\Theta_{s,t}$, $s\leq t$ satisfies the $p$-Foster-Lyapunov condition, and let $\overline{b},\overline{C},h>0$ be such that (3.2) holds. Let $(t_{i})_{i\in{\mathbb{N}}}$ be an increasing sequence such that $t_{0}\geq 0$, $t_{1}-t_{0}\leq h$ and $(t_{i}-t_{i-1})_{i\in{\mathbb{N}}^{\ast}}$ is nonincreasing. Then, for every $n\in{\mathbb{N}}$
@@
$M_{p}(\Theta,\mu)\leq\left\|\mu\right\|_{p}^{p}+\overline{C}\frac{e^{\overline{b}}}{\overline{b}}.$ (3.6)
@@
In many situations to follow in the examples, one applies a sequence of arguments which are a recall of the results of Theorems 2.15, 2.12 and Proposition 2.11.
@@
$W_{p}^{p}(\theta_{s,t}(\mu),\nu)\leq Ae^{-\varepsilon\wedge b_{\ast}(t-s)}(1+(t-s)^{\mathbf{1}_{b_{\ast}=\varepsilon}})$ (3.8)
with $A$ being as given by Corollary 2.9.
Furthermore, let $\gamma_{n}=\frac{1}{n+h^{-1}}$ and $t_{n}=s+\gamma_{1}+...+\gamma_{n}$, $n\in{\mathbb{N}}$. Then, we have for every $\mu\in\mathcal{P}_{p}(B)$
with $A=2^{2p}C_{*}(1+M_{p}(\theta,\mu))(2C_{b_{\ast},\varepsilon}+1)$.
@@ 4. Examples with non global contraction conditions (Langevin type equations)
In several recent papers (see Eberle [Ebe16], Schuh [Sch24] ) authors have used a special coupling procedure, called “reflecting coupling” in order to prove convergence to equilibrium for the Langevin equation. An important point is that the result holds under a weak contraction hypothesis (i.e. out of a compact set), in contrast with older results in which uniform construction properties are needed. It turns out that we are able to use Eberle’s result concerning the flow in order to prove convergence of the Euler scheme as well and to control the speed of convergence. First, we consider the classical Langevin equation given by
@@
Suppose that (4.2) and (4.3) hold and let $\theta_{s,t}(\mu)$ be the law of $X_{t}$ when $X_{s}\sim\mu$ is independent of $(B_{u}-B_{s},u\geq s)$. Then, there is a constant $b_{*}>0$ such that
$W_{1}(\theta_{s,t}(\mu),\theta_{s,t}(\nu))\leq W_{1}(\mu,\nu)e^{-b_{\ast}(t-s)}.$ (4.5)
Moreover, $\theta$ has a unique invariant measure $\nu$ which has finite moment of order one.
@@
$\mathbb{E}\left[\left|X_{s,t}-Y_{s,t}\right|\right]\leq\int_{s}^{t}C(1+\mathbb{E}\left|Y\right|)(r-s)^{1/2}dr\leq C(1+\mathbb{E}\left|Y\right|)(t-s)^{3/2}.$
@@
$dY_{t}$ $=$ $(ub^{E}(X_{t})+u\int_{{\mathbb{R}}^{d}}b^{I}(X_{t},z)\mu_{t}^{X}(dz)-\gamma Y_{t})dt+\sqrt{2\gamma u}dB_{t}$
@@
$W_{1}(\mathcal{L}(X_{t},Y_{t}),\mathcal{L}(X_{t}^{\prime},Y_{t}^{\prime})))\leq e^{-b_{\ast}t}W_{1}(\mu,\mu^{\prime}).$ (4.9)
@@
$W_{1}^{|\cdot|}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq Cn^{-\frac{1}{2}\wedge b_{\ast}}(1+\ln(n+1)^{1_{\{b_{\ast}=\frac{1}{2}\}}})$ (4.10)
@@
Let $(X_{t},Y_{t})$ be the solution starting from $(\overline{X},\overline{Y})$ such that $\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)<\infty$. We claim that there exists a constant $C\in{\mathbb{R}}_{+}$ such that for $s\leq t\leq s+1$,
$\mathbb{E}\left|X_{s,t}-\overline{X}_{s,t}\right|+\mathbb{E}\left|Y_{s,t}-\overline{Y}_{s,t}\right|\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right)(t-s)^{3/2}.$ (4.11)
@@
$\mathbb{E}\left|X_{s,t}\right|+\mathbb{E}\left|Y_{s,t}\right|\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right).$ (4.12)
@@
$\mathbb{E}|X_{s,t}-\overline{X}|\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right)(t-s),$
$\mathbb{E}|Y_{s,t}-\overline{Y}|\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right)(t-s)+C(t-s)^{1/2}.$
@@
$\mathbb{E}(\left|X_{s,t}-\overline{X}_{s,t}\right|)\leq\int_{s}^{t}\mathbb{E}(\left|Y_{s,r}-\overline{Y}\right|)dr\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right)(t-s)^{2}+C(t-s)^{3/2},$
@@
$\mathbb{E}(\left|Y_{s,t}-\overline{Y}_{s,t}\right|)\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right)(t-s)^{2}+C(t-s)^{3/2}.$
@@
$\mathbb{E}(d(X_{s,t},\overline{X}_{s,t}))+\mathbb{E}(d(Y_{s,t},\overline{Y}_{s,t}))\leq C\left(1+\mathbb{E}(\left|\overline{X}\right|+\left|\overline{Y}\right|)\right)(t-s)^{3/2},$
@@ 5. Examples with global contraction conditions
In this section, we propose to investigate different examples of stochastic systems that fall into the framework presented in Section 2. In contrast with Section 4 where the coupling property between the approximation scheme and the flow is deduced from the self-coupling property of the flow, we instead present in these section cases where the coupling property is deduced from the self-coupling property of the approximation scheme. When applying it we will need to prove the main coupling properties as described in Definition 2.3, as well as the $p$-boundedness stated in (2.4). In most examples, the main argument is based on an application of Theorem 3.7. Therefore the proofs consist in the verification of (3.4) and (2.22). The first example on reflected SDEs considers a rather simple example with Markovian dynamics without law dependence on the coefficients. Then, we consider a neuronal model, which is an example of McKean-Vlasov SDE. To show that the general framework in Section 2 can be applied in Wasserstein spaces with different value of $p$, we will obtain conditions which imply the convergence of the approximation schemes to the invariant measure in $W_{2}$ and $W_{1}$. This will lead to different contraction hypothesis. In these two examples, we can apply directly the method presented in Subsections 2.3 and 3.2. Namely, we construct a coupling between the continuous stochastic process and its approximation scheme. In contrast, for the Boltzmann equation approximation in $W_{2}$, we are not able to construct a coupling between the Boltzmann process and its one-step approximation scheme. To get around this difficulty, we will couple an approximation of the Boltzmann process (with say parameter $m$) with the one-step approximation scheme and then take limits as $m\to\infty$ in all the estimates. For the Boltzmann equation in $W_{1}$, we can skip this step relying on the sewing lemma. Note that in all the examples, we have not strived to find the best conditions as our goal is to show in a compact form that the main framework applies in each situation. Considering weaker conditions or more general equations in each example may be possible, but outside the scope of the paper. For readability, we have selected some proofs to keep them in the main text while the other ones are available in Appendix B.
@@ 5.1. A stochastic equation with reflection in $W_{2}$
The proofs are based on upper bounds for the $L^{2}$ norms of $\mathcal{X}^{1}(X^{1})-\mathcal{X}^{2}(X^{1})$ as well as $\mathcal{X}^{1}(X^{1})-\mathcal{X}^{1}(X^{2})$. These are performed using Ito’s formula and then the hypotheses give that the joint probabilistic representation is satisfied as stated in Definition 2.14. We refer to Appendix B for the details.
5.2. A neuronal model
@@
We assume moreover the following contraction condition for $x,y\in\mathbb{R}_{+}$
$2(y-x)(b(y)-b(x))\leq-\bar{b}(y-x)^{2}+y^{2}f_{M}(y)+x^{2}f_{M}(x)-2xyf_{M}(y)\vee f_{M}(x).$ (5.9)
This condition combines drift and jump effects in the model to create the contraction.
@@
Let us note that $\mathcal{X}^{2}_{s,t}$ stands for the approximation of the continuous process $\mathcal{X}^{1}_{s,t}$ but is different from the Euler scheme. Indeed, the coefficient of the driving Poisson measure depends on $\mathcal{X}^{2}_{s,u-}(X^{2})$, not on $X_{2}$. This plays a crucial role in the analysis below, in particular for the term (5.14). Nonetheless, the process $\mathcal{X}^{2}_{s,t}(X^{2})$ can be sampled exactly, as a piecewise deterministic Markov process.

### v2

@@ 2.3. Probabilistic representation
$W_{p}^{p}(\mu^{1},\mu^{2})=$ $\mathbb{E}[d^{p}(X^{1},X^{2})],$ (2.26)
$\mathbb{E}[d^{p}(\mathcal{X}_{s,t}^{1},\mathcal{X}_{s,t}^{2})]\leq$ $(1-b_{\ast}(t-s))\mathbb{E}[d^{p}(X^{1},X^{2})]+C_{\ast}\Gamma_{p}(\mu^{1},\mu^{2})(t-s)^{1+\varepsilon}.$ (2.27)
@@
$W_{p}^{p}(\Theta_{s,t}^{1}(\mu^{1}),\Theta_{s,t}^{2}(\mu^{2}))$ $\leq\mathbb{E}[d^{p}(\mathcal{X}_{s,t}^{1},\mathcal{X}_{s,t}^{2})]$
$\leq(1-b_{\ast}(t-s))\mathbb{E}[d^{p}(X^{1},X^{2})]+C_{\ast}\Gamma_{p}(\mu^{1},\mu^{2})(t-s)^{1+\varepsilon}$
@@
$M_{p}(\Theta,\mu,h)\leq\left\|\mu\right\|_{p}^{p}+C\frac{e^{b}}{b}.$ (2.28)
@@ 3.1. Using the sewing lemma to prove coupling
We thus get that $\Theta^{1}$ and $\Theta^{2}$ are $(1,b_{\ast},\varepsilon)$-coupled. We then use Lemma 2.6 and pick $h^{\prime}\in(0,h)$ that fulfills (2.11) to get that $\Theta^{2}$ is also $1$-bounded. ∎
@@
  • If we are able to prove independently (relying on continuous time arguments) that $\theta$ is $1$-bounded and is $(1,b_{\ast},\varepsilon)$ self-coupled, then $\Theta$ and $\theta$ are $(1,b_{\ast},\varepsilon)$-coupled and $\Theta$ is $1$-bounded. In particular, if $\theta$ is time homogeneous and $\nu$ is the invariant measure for $\theta$, and if we take the time sequence $(t_{n})$ as in Theorem 2.15 then, by (2.23), for every $\mu\in\mathcal{P}_{1}(B)$,
@@ 3.2. A Foster-Lyapunov type criterion to prove $p$-boundedness
Suppose that the family $\Theta_{s,t}$, $s\leq t$ satisfies the $p$-Foster-Lyapunov condition, and let $\overline{b},\overline{C},h>0$ be such that (3.2) holds. Let $(t_{i})_{i\in{\mathbb{N}}}$ be an increasing sequence such that $t_{0}\geq 0$ and $t_{i}-t_{i-1}\leq h$ for $i\in{\mathbb{N}}^{\ast}$. Then, for every $n\in{\mathbb{N}}$
@@
$M_{p}(\Theta,\mu{,h})\leq\left\|\mu\right\|_{p}^{p}+\overline{C}\frac{e^{\overline{b}}}{\overline{b}}.$ (3.6)
@@
In many situations to follow in the examples, one applies a sequence of arguments which are a recall of the results of Proposition 2.19, Theorem 2.15 and Proposition 2.14. They are summarized in the next theorem.
@@
$W_{p}^{p}(\theta_{s,t}(\mu),\nu)\leq C(1+\|\mu\|_{p}^{p})e^{-\varepsilon\wedge b_{\ast}(t-s)}(1+(t-s)^{\mathbf{1}_{b_{\ast}=\varepsilon}})$ (3.8)
where the constant $C\in{\mathbb{R}}_{+}$ does not depend on $\mu$.
Furthermore, let $a\geq 0$, $\gamma_{n}=\frac{1}{a+n}$ and $t_{n}=s+\gamma_{1}+...+\gamma_{n}$, $n\in{\mathbb{N}}$. Then, there exists $C\in{\mathbb{R}}_{+}$ such that for every $\mu\in\mathcal{P}_{p}(B)$,
$W_{p}^{p}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq C(1+\|\mu\|_{p}^{p})n^{-b_{\ast}\wedge\varepsilon}(1+\ln(n+1)^{\mathbf{1}_{b_{\ast}=\varepsilon}}).$ (3.9)
In contrast with Theorem 2.4, we assume here that the Foster Lyapunov criterion (3.4) hold for $\theta$ and $\Theta$, which gives a uniform estimate with respect to $\mu$, see Remark 2.16.
@@ 4. Examples with non global contraction conditions (Langevin type equations)
In several recent papers (see in particular Eberle [Ebe16], Eberle et al. [EGZ19] and Schuh [Sch24] ) authors have used a special coupling procedure, called “reflecting coupling” in order to prove convergence to equilibrium for the Langevin type equations. An important point is that the result holds under a weak contraction hypothesis (i.e. out of a compact set), in contrast with older results in which uniform construction properties are needed. It turns out that we are able to use Eberle’s result concerning the flow in order to prove convergence of the Euler scheme as well and to control the speed of convergence. First, we consider the classical Langevin equation given by
@@
Suppose that (4.2) and (4.3) hold and let $\theta_{s,t}(\mu)$ be the law of $X_{t}$ when $X_{s}\sim\mu$ is independent of $(B_{u}-B_{s},u\geq s)$. Then, there is a constant $b_{*}>0$ such that for any $\mu^{1},\mu^{2}\in\mathcal{P}_{1}({\mathbb{R}}^{d})$,
$W_{1}(\theta_{s,t}(\mu^{1}),\theta_{s,t}(\mu^{2}))\leq W_{1}(\mu^{1},\mu^{2})e^{-b_{\ast}(t-s)}.$ (4.5)
Moreover, $\theta$ has a unique invariant measure $\nu\in\mathcal{P}_{1}({\mathbb{R}}^{d})$.
@@
$\mathbb{E}\left[\left|X_{s,t}-Y_{s,t}\right|\right]\leq\int_{s}^{t}C(1+\mathbb{E}[\left|Y\right|])(r-s)^{1/2}dr\leq C(1+\mathbb{E}[\left|Y\right|])(t-s)^{3/2}.$
@@
$dY_{t}$ $=(ub^{E}(X_{t})+u\int_{{\mathbb{R}}^{d}}b^{I}(X_{t},z)\mu_{t}^{X}(dz)-\gamma Y_{t})dt+\sqrt{2\gamma u}dB_{t},$
@@
$W_{1}(\mathcal{L}(X_{t},Y_{t}),\mathcal{L}(X_{t}^{\prime},Y_{t}^{\prime}))\leq e^{-b_{\ast}t}W_{1}(\mu,\mu^{\prime}).$ (4.9)
@@
$W_{1}^{|\cdot|}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq Cn^{-\frac{1}{2}\wedge b_{\ast}}(1+\ln(n+1)^{1_{\{b_{\ast}=\frac{1}{2}\}}}).$ (4.10)
@@
Let $(X_{t},Y_{t})$ be the solution starting from $(\overline{X},\overline{Y})$ such that $\mathbb{E}(|\overline{X}|+|\overline{Y}|)<\infty$. We claim that there exists a constant $C\in{\mathbb{R}}_{+}$ such that for $s\leq t\leq s+1$,
$\mathbb{E}[|X_{s,t}-\overline{X}_{s,t}|]+\mathbb{E}[|Y_{s,t}-\overline{Y}_{s,t}|]\leq C\left(1+\mathbb{E}[|\overline{X}|+|\overline{Y}|]\right)(t-s)^{3/2}.$ (4.11)
@@
$\mathbb{E}[|X_{s,t}|]+\mathbb{E}[|Y_{s,t}|]\leq C\left(1+\mathbb{E}[|\overline{X}|+|\overline{Y}|]\right).$ (4.12)
@@
$\mathbb{E}[|X_{s,t}-\overline{X}|]\leq C\left(1+\mathbb{E}[|\overline{X}|+|\overline{Y}|]\right)(t-s),$
$\mathbb{E}[|Y_{s,t}-\overline{Y}|]\leq C\left(1+\mathbb{E}[|\overline{X}|+|\overline{Y}|]\right)(t-s)+C(t-s)^{1/2}.$
@@
$\mathbb{E}[|X_{s,t}-\overline{X}_{s,t}|]\leq\int_{s}^{t}\mathbb{E}[|Y_{s,r}-\overline{Y}|]dr\leq C\left(1+\mathbb{E}[|\overline{X}|+|\overline{Y}|]\right)(t-s)^{2}+C(t-s)^{3/2},$
@@
$\mathbb{E}[|Y_{s,t}-\overline{Y}_{s,t}|]\leq C\left(1+\mathbb{E}[|\overline{X}|+|\overline{Y}|]\right)(t-s)^{2}+C(t-s)^{3/2}.$
@@
$\mathbb{E}[d(X_{s,t},\overline{X}_{s,t})]+\mathbb{E}[d(Y_{s,t},\overline{Y}_{s,t})]\leq C\left(1+\mathbb{E}[|\overline{X}|+|\overline{Y}|]\right)(t-s)^{3/2},$
@@ 5. Examples with global contraction conditions
In this section, we propose to investigate different examples of stochastic systems that fall into the framework presented in Section 2. In contrast with Section 4 where the coupling property between the approximation scheme and the flow is deduced from the self-coupling property of the flow, we instead present in these section cases where the coupling property is deduced from the self-coupling property of the approximation scheme. When applying it we will need to prove the main coupling properties as described in Definition 2.2, as well as the $p$-boundedness stated in Definition 2.3. In most examples, the main argument is based on an application of Theorem 3.7. Therefore the proofs consist in the verification of (3.4) and (2.27). The first example on reflected SDEs considers a rather simple example with Markovian dynamics without law dependence on the coefficients. Then, we consider a neuronal model, which is an example of McKean-Vlasov SDE. To show that the general framework in Section 2 can be applied in Wasserstein spaces with different value of $p$, we will obtain conditions which imply the convergence of the approximation schemes to the invariant measure in $W_{2}$ and $W_{1}$. This will lead to different contraction hypothesis. In these two examples, we can apply directly the method presented in Subsections 2.3 and 3.2. Namely, we construct a coupling between the continuous stochastic process and its approximation scheme. In contrast, for the Boltzmann equation approximation in $W_{2}$, we are not able to construct a coupling between the Boltzmann process and its one-step approximation scheme. To get around this difficulty, we will couple an approximation of the Boltzmann process (with say parameter $m$) with the one-step approximation scheme and then take limits as $m\to\infty$ in all the estimates. For the Boltzmann equation in $W_{1}$, we can skip this step relying on the sewing lemma. Note that in all the examples, we have not strived to find the best conditions as our goal is to show in a compact form that the main framework applies in each situation. Considering weaker conditions or more general equations in each example may be possible, but outside the scope of the paper. For readability, we have selected some proofs to keep them in the main text while the other ones are available in Appendix C.
@@ 5.1. A stochastic equation with reflection in $W_{2}$
The proofs are based on upper bounds for the $L^{2}$ norms of $\mathcal{X}^{1}(X^{1})-\mathcal{X}^{2}(X^{1})$ as well as $\mathcal{X}^{1}(X^{1})-\mathcal{X}^{1}(X^{2})$. These are performed using Ito’s formula and then the hypotheses give that the joint probabilistic representation is satisfied as stated in Definition 2.18. We refer to Appendix C for the details.
5.2. A neuronal model in $W_{1}$
@@
We assume moreover the following contraction condition: there exists $\bar{b}>0$ such that for any $x,y\geq 0$,
$\mathrm{sgn}(y-x)(b(y)-b(x))-|y-x|(f_{M}(y)\vee f_{M}(x))$
$+x(f_{M}(y)-f_{M}(x))^{+}+y(f_{M}(x)-f_{M}(y))^{+}\leq-\bar{b}\left|x-y\right|.$ (5.9)
This condition combines drift and jump effects in the model to create the contraction. We however remark that in the particular case that $f_{M}$ and ${\mathbb{R}}_{+}\ni x\mapsto\frac{x}{f_{M}(x)}$ are nondecreasing functions (which holds for example if $f(x)=x^{a}$ with $a\in[0,1]$), we have
$-|y-x|(f_{M}(y)\vee f_{M}(x))+x(f_{M}(y)-f_{M}(x))^{+}+y(f_{M}(x)-f_{M}(y))^{+}\leq 0,$
since, for $x<y$ we have $(y-x)f_{M}(y)+x(f_{M}(y)-f_{M}(x))\leq-yf_{M}(y)+xf_{M}(y)+yf_{M}(x)-xf_{M}(x)=-(y-x)(f_{M}(y)-f_{M}(x))\leq 0$. In this case, (5.9) is satisfied if $\mathrm{sgn}(y-x)(b(y)-b(x))\leq-\bar{b}\left|x-y\right|$.
@@
Let us note that $\mathcal{X}^{2}_{s,t}$ stands for the approximation of the continuous process $\mathcal{X}^{1}_{s,t}$ but is different from the Euler scheme. Indeed, the coefficient of the driving Poisson measure depends on $\mathcal{X}^{2}_{s,u-}(X^{2})$, not on $X_{2}$. This plays a crucial role in the analysis below, precisely for the term named $I_{3}(s,t)$ in the proof of Theorem 5.3. Nonetheless, the process $\mathcal{X}^{2}_{s,t}(X^{2})$ can be sampled exactly, as a piecewise deterministic Markov process.